Suppose that on January 1 you have a balance of ​$3700 on a credit card whose APR is 15​%, which you want to pay off in 1 year.

Monthly payment?

Total paid?

The percentage of the total paid that is interest is?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

BC $3,700.00 15% Credit card balance APR Monthly interest rate No. of monthly payments 1.25% 12 Amount of Monthly payment $33

Cell reference -

3700 0.15 Credit card balance APR Monthly interest rate No. of monthly payments 4 =C3/12 12 =PMT(C4,C5,-C2,,0) 7 a. Amount of

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.